ACID-BASE ELECTRONIC PROPERTIES IN THE GAS PHASE: PERMANENT ELECTRIC DIPOLE MO-MENTS OF A PHOTOACIDIC SUBSTRATE.^{a}

\underline{ADAM J. FLEISHER}, PHILIP J. MORGAN and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, 15260.

The permanent electric dipole moments of two conformers of 2-naphthol (2HN) in their ground and electronically excited states have been experimentally determined by Stark-effect measurements in a molecular beam. When in solution, 2HN is a weak base in the S_{0} state and a strong acid in the S_{1} state.^{b} Using sequential solvation of the cis-2HN photoacid with the base ammonia, we have begun to approach condensed phase acid-base interactions with gas phase rotational resolution.^{c} Our study, void of bulk solvent perturbations, is of importance to the larger community currently describing aromatic biomolecule and ”super” photoacid behavior via theoretical modeling and condensed phase solvatochromism.

aWork supported by NSF (CHE-0615755).

bA. Weller. Prog. React. Kinet. 5, 273 (1970).

cD. F. Plusquellic, X. -Q. Tan, and D. W. Pratt. J. Chem. Phys. 96, 8026 (1992).